OpenStudy (anonymous):

stumped...trying to figure out the y intercept for -3x^2-6x-5

OpenStudy (amistre64):

well, since the y intercept is defined when x=0 .....

OpenStudy (amistre64):

zero out all the x parts and what left over?

OpenStudy (anonymous):

i got 5/3=x?

OpenStudy (amistre64):

how? multiplying by 0 is one of the easier things to do in math

OpenStudy (amistre64):



OpenStudy (amistre64):

y = -3x^2-6x-5 ; now when x = 0 we gots.. y = -3(0)^2- 6(0) -5 y = 0 -0-5 ...

OpenStudy (anonymous):

i got that for y but i'm trying to solve x

OpenStudy (amistre64):

lol .... the Y INTERCEPT is defined when: x = 0

OpenStudy (anonymous):

(0,5) is my Y

OpenStudy (amistre64):

that is it

OpenStudy (anonymous):

okay so my X intercept will be (0,0)?

OpenStudy (amistre64):

no, thats a whole different can o worms

OpenStudy (anonymous):

just what i was asking for.

OpenStudy (aravindg):

amistre help me

OpenStudy (amistre64):

X intercept is defined when y = 0 sooo lets factor; if we can y = -3x^2 -6x -5 = 0 -(3x^2 +6x +5) = 0 -3(x+5/3)(x+1/3) = 0 when either (x+ 5/3) = 0 or (x+1/3) = 0

OpenStudy (amistre64):

x = -5/3 or x = -1/3 as i see it

OpenStudy (anonymous):

right! i got that earlier.

OpenStudy (amistre64):

good :)

OpenStudy (anonymous):

so now i divide -5/3 to plot on graph?

OpenStudy (anonymous):

X=1.6

OpenStudy (amistre64):

divide? no somehow i jumped the gun

OpenStudy (amistre64):

there are no x intercepts since the graph never touches it

OpenStudy (anonymous):

so i can't plug in 0 for y and get the x intercept?

OpenStudy (anonymous):

hello